alextrejo373:

Based on your data, what is the experimental probability that the family has two dogs or two cats? If the family has three pets, what is the theoretical probability that they have three dogs or three cats? How could you change the simulation to generate data for three pets?

alextrejo373:

Here is a Picture

Unnecessary:

Is there more to the assignment?

alextrejo373:

No I just need help with the questions at the bottom of the picture

Unnecessary:

Also I noticed that in question 1) you stated that the theoretical probability that a family has two dogs or two cats is 25/50 or 1/2 then you wrote "or it also 25%" Not sure what you mean by this because 25/50 is 1/2 but that would be 50% not 25%.

alextrejo373:

ok

Unnecessary:

Also it seems like you're leaving something out, you either need to have data on the families that have either two cats or two dogs in order to answer the questions. Is there not a table or text with this information?

Unnecessary:

Oooooh, I think I get it now.

Unnecessary:

The coin flip data represents whether or not it was a cat or not.

Unnecessary:

I guess you can just assume that heads = dog, and tails = cats.

Unnecessary:

It says that the probability of getting two cats would be 13/50 or 26%, and the probabiltiy of getting two dogs would be 14/50 or 28%.

Extrinix:

@unnecessary wrote:
It says that the probability of getting two cats would be 13/50 or 26%, and the probabiltiy of getting two dogs would be 14/50 or 28%.
It says the probability of 2 cats `or` 2 dogs, making it \(\dfrac{13+14}{50}\) = \(\dfrac{27}{50}\), or \(\dfrac{54}{100}\) So they would have a 54% / 0.54 chance of getting 2 dogs and 2 cats.